OpenStudy (anonymous):

simplify the radical expression sqrt72x^2 ?

OpenStudy (kc_kennylau):

\[\sqrt{72}=\sqrt{2^3*3^2}=2*3*\sqrt2=6\sqrt2\]

OpenStudy (kc_kennylau):

Is the x^2 inside or outside?

OpenStudy (anonymous):

inside. I don't know how to make the square root sign